[Brahms-l] Shift report 20050517 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue May 17 2005 - 08:00:41 EDT
Shifter leader was Thomas Materna
Shifters were 


1. Beam during ~6h30
2. Data run 14705-14716

3. All above runs have HV problem. 
   Mainframe BB3 had some hardware problem (according to Fleming) 
   and cannot reset the proper HVs for those connected to that 
   mainframe. Those are BBR31-35 BBL37-44, CCL1-19. 
   In addition, H3 Slats has changed slightly from run 1414713 to 14714.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 17 08:01:19 2005

This archive was generated by hypermail 2.1.8 : Tue May 17 2005 - 08:01:27 EDT